OpenStudy (anonymous):

Find the distance between points (-4, -6) and (5, -1). THIS is what i have now: sqroot (-9)^2 + (-5)^2

OpenStudy (anonymous):

I'm confused on what to do next ....

OpenStudy (radar):

Sum those squares and then take the square root of the sum.

OpenStudy (anonymous):

Find the distance between points (-4, -6) and (5, -1).

OpenStudy (anonymous):

I did the second part..

OpenStudy (radar):

\[\sqrt{106}\]

OpenStudy (anonymous):

-9^2 = invalid?

OpenStudy (radar):

(-9)(-9)=81

OpenStudy (radar):

No not invalid, you have the correct data, just use it.

OpenStudy (anonymous):

oh okay. thank you !

OpenStudy (radar):

You are using the distance formula which is\[\sqrt{(x1-x2)^{2}+(y1-y2)^{2}}\]\[\sqrt{(-9)^{2}+(-5)^{2}}=\sqrt{106}\]

OpenStudy (anonymous):

can you help me with this? . What is the equation of a parabola with vertex (0, 0) and directrix x = -3. Write the equation of an ellipse centered at the origin with foci at (0, -3) and (0, 3) and a major axis of 10.

OpenStudy (anonymous):

those are my last qeustions

OpenStudy (radar):

I need to review these for a while, the term directrix is unfamiliar to me, I know very little about it sorry, repost the question and I believe there are many that will help who are familiar with directrix and ellipse.